# Logarithmic delocalization of low temperature 3D Ising and Potts interfaces above a hard floor

### v1

@@ 2.4.1. Map to lift $\mathcal{I}_{\mathsf{top}}$
Let $\omega_{0}$ be the configuration such that $(\omega_{0})_{e}=0$ if and only if $e$ is dual to a plaquette $f\in I_{0}$. It suffices to show that for any $\omega$ with $\mathcal{I}_{\mathsf{full}}(\omega)=I_{0}$, we have $\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)=\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})$, as that would imply $\mathcal{I}_{\mathsf{top}}(\omega)=\mathcal{I}_{\mathsf{top}}(\omega_{0})$. Since $\omega$ must be pointwise below $\omega_{0}$, we have immediately that $\mathcal{V}_{\mathsf{top}}(\omega)\subset\mathcal{V}_{\mathsf{top}}(\omega_{0})$, and hence $\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)\subset\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})$. To show the other inclusion, suppose for contradiction that $\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})\setminus\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)$ is nonempty. Then, there must in particular be some $v\in\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})\setminus\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)$ such that $v$ is adjacent to some $u\in\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})^{c}$ (if no such $v$ exists, then $\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})\setminus\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)$ must consist of finite components surrounded by vertices of $\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})\cap\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)$, but this contradicts the construction of $\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)$ which would have included such finite components). Let $C(v)$ be the finite random cluster component containing $v$. Then $C(v)$ has a 1-connected boundary of dual-to-closed plaquettes which contains the plaquette dual to the edge $(u,v)$, hence the entire 1-connected boundary is a subset of $\mathcal{I}_{\mathsf{full}}$. Thus, $v$ is in a finite component even in $\omega_{0}$, which together with the fact that $v$ is adjacent to $u\in\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})^{c}$ implies that $v\in\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})^{c}$, concluding the contradiction. ∎
@@
We claim that it will therefore suffice to show that there is a variant of the lifting map $\Theta_{j}$ for $\mathsf{full}$ interfaces, call it $\Phi_{j}$, such that for every $I$ and every $I_{0}\in\mathcal{F}(I)$, we have
@@
For every $j\geq 1$ and every realization $I_{0}$ for $\mathcal{I}_{\mathsf{full}}$ under $\bar{\pi}_{n}^{\mathsf{dob}}$, the map $\Phi_{1}$ outputs a valid full interface $J_{0}=\Phi_{1}(I_{0})$ separating $\partial_{\mathsf{blue}}\Lambda^{\prime}_{n,n}$ from $\partial_{\mathsf{red}}\Lambda^{\prime}_{n,n}$.
@@
To show that $J_{0}$ is separating, suppose for contradiction that there is a path $P$ from $x\in\partial_{\mathsf{red}}\Lambda^{\prime}_{n,n}$ to $y\in\partial_{\mathsf{blue}}\Lambda^{\prime}_{n,n}$ that does not cross any plaquette of $J_{0}$. Let us first consider the case where $x\in\mathcal{L}_{\geq 3/2}$. Then we can consider the translation $P^{\prime}$ of the path $P$ by height $-1$. Since $P$ does not cross any plaquette of $J_{0}$, the path $P^{\prime}$ does not cross any plaquette of $I_{0}$; so, the fact $x-{\mathfrak{e}_{3}}\in\partial_{\mathsf{red}}\Lambda^{\prime}_{n,n}$ implies that $P^{\prime}$ is a path from $\partial_{\mathsf{red}}\Lambda^{\prime}_{n,n}$ to $\partial_{\mathsf{blue}}\Lambda^{\prime}_{n,n}$ that does not cross a plaquette of $I_{0}$, contradicting the fact that $I_{0}$ is a valid interface. For the remaining case $x\in\mathcal{L}_{1/2}\cap\partial\Lambda^{\prime}_{n,n}=:X$, the new (rather than shifted) plaquettes added to $J_{0}$ (of which there are at most $4n$) ensure that any path in $\Lambda^{\prime}_{n,n}$ from $X$ to $X^{c}$ must first pass through a vertex $x^{\prime}\in\mathcal{L}_{3/2}\cap\partial\Lambda^{\prime}_{n,n}$, putting us back in the first case. ∎
@@
The first term is bounded by $4Kn$, and we can bound the second term since the radius $r(f,I_{0};f^{\prime},J_{0})$ is obtained by the distance to either one of the up to $4n$ plaquettes in $J_{0}\setminus I_{1}$ or one of the $4n$ horizontal boundary plaquettes at height 00 in the Dobrushin boundary conditions. If the union of these two plaquette sets is called $F$, then we have
@@ 2.4.2. Injecting entropy via downward spikes
Fix any $I\in\mathbf{I}_{\mathsf{bad}}$. Consider the operation $\Theta_{J}$ which lifts $I$ by $H$ and adds all $4Hn$ vertical plaquettes along the boundary at heights between 00 and $H$. By Lemma 2.15 and Lemma 2.20, we have
@@
    closing the $4h$ horizontal edges (dual to vertical plaquettes) incident $(v_{i}^{x})_{i=1,\ldots,h}$, as well as the one vertical edge (dual to a horizontal plaquette) between $v_{h}^{x}$ and $v_{h}^{x}-{\mathfrak{e}_{3}}$ (if they are not closed already).
@@
11 11 $\geq\sum_{I\in\mathbf{I}_{\mathsf{bad}}}\,\sum_{A\subset\{x:\underline{\operatorname{ht}}_{x}I\leq h\}}\,\sum_{\omega^{\prime}\in\Phi_{A}(\{\omega:\mathcal{I}_{\mathsf{top}}(\omega)=\Theta_{H}I\})}\bar{\pi}_{n}^{\mathsf{fl}}(\omega^{\prime})$
@@
where the right-hand side of (2.30) became $e^{-(4\beta+C)h|A|}$. Using that for every $I\in\mathbf{I}_{\mathsf{bad}}$, one has |$|\{x:\underline{\operatorname{ht}}_{x}I\leq h\}|>\varepsilon_{\beta}n^{2}$, then up to a change of the constant $C$, one has
11 11 $\geq\sum_{I\in\mathbf{I}_{\mathsf{bad}}}\bar{\pi}_{n}^{\mathsf{fl}}(\mathcal{I}_{\mathsf{top}}=\Theta_{H}I)\Big{(}1+e^{-(4\beta+C)h}\Big{)}^{|\{x:\underline{\operatorname{ht}}_{x}I\leq h\}|}\geq\sum_{I\in\mathbf{I}_{\mathsf{bad}}}\bar{\pi}_{n}^{\mathsf{fl}}(\mathcal{I}_{\mathsf{top}}=\Theta_{H}I)\Big{(}1+e^{-(4\beta+C)h}\Big{)}^{\varepsilon_{\beta}n^{2}}\,.$
@@
11 11 $\geq\Big{(}1+e^{-(4\beta+C)h}\Big{)}^{\varepsilon_{\beta}n^{2}}\sum_{I\in\mathbf{I}_{\mathsf{bad}}}e^{-(4\beta+C)Hn}\bar{\pi}_{n}^{\mathsf{fl}}(\mathcal{I}_{\mathsf{top}}=I)$
@@ 2.5. Extending the lower bound to the blue interface
We would now like to extend to $\mathcal{I}_{\mathsf{blue}}$ the lower bound of $\frac{1}{4\beta+C}\log n$ obtained in Proposition 2.19 on the height of most plaquettes in $\mathcal{I}_{\mathsf{top}}$. Note that, to that end, the additive error of $(8/\beta)\log n$ from Corollary 2.14 (whereby $\mathcal{I}_{\mathsf{blue}}$ is sandwiched between $\mathcal{I}_{\mathsf{bot}},\mathcal{I}_{\mathsf{top}}$) would be far too crude; instead, we will rely on the following result.
@@
with $K>0$ the absolute constant from Proposition 2.3 and $C>0$ another absolute constant. Note that, unlike the more careful analysis of the map $\Phi_{j}$ in Section 2.4.1 (which would correspond to bounding $|\sum_{f\in I}{\mathbf{g}}(f,I)-\sum_{f^{\prime}\in I_{0}}{\mathbf{g}}(f^{\prime},I_{0})|$ via absorbing $K$ for each $f\in I\oplus I_{0}$, and another constant contribution for each such $f$ from summing over the exponential decay of its effect on other plaquettes through the ${\mathbf{g}}$ term) we resorted to the naive bound $(|I|+|I_{0}|)\|{\mathbf{g}}\|_{\infty}\leq(2+s)Kn^{2}$; this is due to $I_{0}$ being adjacent to the boundary of the box, whereby $r(f,I;f^{\prime},I_{0})$ is $1$ for all $f^{\prime}\in I_{0}$ (demonstrating some of the difficulty in overcoming the potential pinning effect of the hard floor).
@@ 3. Sharp upper bound
    there exists a separating set $\mathcal{S}_{N}$ in $\omega\wedge\omega^{\prime}$ between $\mathcal{L}_{3\ell/4}$ and $\mathcal{L}_{\ell}$
@@
On the event $\mathcal{G}$, expose the highest realizing $\mathcal{S}_{N}$ and the lowest realizing $\mathcal{S}_{S}$. By the domain Markov property, we can resample the configurations $(\omega,\sigma)$ and $(\omega^{\prime},\sigma^{\prime})$ in the region sandwiched between these separating surfaces using the identity coupling (the colorings on $\mathcal{S}_{N}$ must be $\mathsf{red}$ in both, and on $\mathcal{S}_{S}$ $\mathsf{blue}$ in both, so indeed the measures induced on the regions sandwiched by the separating surfaces have the same distribution). Since $\mathcal{I}_{\mathsf{full}}$ and therefore $\mathcal{I}_{\mathsf{blue}}$ must be sandwiched between these two separating surfaces when $|\mathsf{h}|\leq\ell/2$, on the event $\mathcal{G}$ we have
@@
The argument for the floor measures $\mu_{\Lambda_{n,\ell}}^{\mathsf{fl}}$ and $\mu_{\Lambda_{n,m}}^{\mathsf{fl}}$ is analogous, only needing the existence of $\mathcal{S}_{N}$, and no need for $\mathcal{S}_{S}$ since the southern boundary is already shared at height zero. ∎

### v2

@@ 2.4.1. Map to lift $\mathcal{I}_{\mathsf{top}}$
Let $\omega_{0}$ be the configuration such that $(\omega_{0})_{e}=0$ if and only if $e$ is dual to a plaquette $f\in I_{0}$. It suffices to show that for any $\omega$ with $\mathcal{I}_{\mathsf{full}}(\omega)=I_{0}$, we have $\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)=\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})$, as that would imply $\mathcal{I}_{\mathsf{top}}(\omega)=\mathcal{I}_{\mathsf{top}}(\omega_{0})$. Since $\omega$ must be pointwise below $\omega_{0}$, we have immediately that $\mathcal{V}_{\mathsf{top}}(\omega)\subset\mathcal{V}_{\mathsf{top}}(\omega_{0})$, and hence $\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)\subset\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})$. To show the other inclusion, suppose for contradiction that $\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})\setminus\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)$ is nonempty. Then, there must in particular be some $v\in\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})\setminus\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)$ such that $v$ is adjacent to some $u\in\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})^{c}$ (if no such $v$ exists, then $\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})\setminus\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)$ must consist of finite components surrounded by vertices of $\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})\cap\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)$, but this contradicts the construction of $\widehat{\mathcal{V}}_{\mathsf{top}}(\omega)$ which would have included such finite components). Let $C(v)$ be the finite random-cluster component containing $v$. Then $C(v)$ has a 1-connected boundary of dual-to-closed plaquettes which contains the plaquette dual to the edge $(u,v)$, hence the entire 1-connected boundary is a subset of $\mathcal{I}_{\mathsf{full}}$. Thus, $v$ is in a finite component even in $\omega_{0}$, which together with the fact that $v$ is adjacent to $u\in\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})^{c}$ implies that $v\in\widehat{\mathcal{V}}_{\mathsf{top}}(\omega_{0})^{c}$, concluding the contradiction. ∎
@@
We claim that it will therefore suffice to show that there is a variant of the lifting map $\Theta_{1}$ for $\mathsf{full}$ interfaces, call it $\Phi_{1}$, such that for every $I$ and every $I_{0}\in\mathcal{F}(I)$, we have
@@
For every realization $I_{0}$ for $\mathcal{I}_{\mathsf{full}}$ under $\bar{\pi}_{n}^{\mathsf{dob}}$, the map $\Phi_{1}$ outputs a valid full interface $J_{0}=\Phi_{1}(I_{0})$ separating $\partial_{\mathsf{blue}}\Lambda^{\prime}_{n,n}$ from $\partial_{\mathsf{red}}\Lambda^{\prime}_{n,n}$.
@@
To show that $J_{0}$ is separating, suppose for contradiction that there is a path $P$ from $x\in\partial_{\mathsf{red}}\Lambda^{\prime}_{n,n}$ to $y\in\partial_{\mathsf{blue}}\Lambda^{\prime}_{n,n}$ that does not cross any plaquette of $J_{0}$. Let us first consider the case where $x\in\mathcal{L}_{\geq 3/2}$. Then we can consider the translation $P^{\prime}$ of the path $P$ by height 1-1. Since $P$ does not cross any plaquette of $J_{0}$, the path $P^{\prime}$ does not cross any plaquette of $I_{0}$; so, the fact $x-{\mathfrak{e}_{3}}\in\partial_{\mathsf{red}}\Lambda^{\prime}_{n,n}$ implies that $P^{\prime}$ is a path from $\partial_{\mathsf{red}}\Lambda^{\prime}_{n,n}$ to $\partial_{\mathsf{blue}}\Lambda^{\prime}_{n,n}$ that does not cross a plaquette of $I_{0}$, contradicting the fact that $I_{0}$ is a valid interface. For the remaining case $x\in\mathcal{L}_{1/2}\cap\partial\Lambda^{\prime}_{n,n}=:X$, the new (rather than shifted) plaquettes added to $J_{0}$ (of which there are at most $4n$) ensure that any path in $\Lambda^{\prime}_{n,n}$ from $X$ to $X^{c}$ must first pass through a vertex $x^{\prime}\in\mathcal{L}_{3/2}\cap\partial\Lambda^{\prime}_{n,n}$, putting us back in the first case. ∎
@@
The first term is bounded by $4Kn$, and we can bound the second term since the radius $r(f,I_{0};f^{\prime},J_{0})$ is obtained by the distance to either one of the up to $4n$ plaquettes in $J_{0}\setminus I_{1}$ or one of the $4n$ horizontal boundary plaquettes at height 0 in the Dobrushin boundary conditions. If the union of these two plaquette sets is called $F$, then we have
@@ 2.4.2. Injecting entropy via downward spikes
Fix any $I\in\mathbf{I}_{\mathsf{bad}}$. Consider the operation $\Theta_{H}$ which lifts $I$ by $H$ and adds all $4Hn$ vertical plaquettes along the boundary at heights between 0 and $H$. By Lemma 2.16 and Lemma 2.21, we have
@@
    closing the $4h$ horizontal edges (dual to vertical plaquettes) incident to vertices $(v_{i}^{x})_{i=1,\ldots,h}$, as well as the one vertical edge (dual to a horizontal plaquette) between $v_{h}^{x}$ and $v_{h}^{x}-{\mathfrak{e}_{3}}$ (if they are not closed already).
@@
$1$ $\geq\sum_{I\in\mathbf{I}_{\mathsf{bad}}}\,\sum_{A\subset\{x:\underline{\operatorname{ht}}_{x}I\leq h\}}\,\sum_{\omega^{\prime}\in\Phi_{A}(\{\omega:\mathcal{I}_{\mathsf{top}}(\omega)=\Theta_{H}I\})}\bar{\pi}_{n}^{\mathsf{fl}}(\omega^{\prime})$
@@
where the right-hand side of (2.30) became $e^{-(4\beta+C)h|A|}$. Using that for every $I\in\mathbf{I}_{\mathsf{bad}}$, one has $|\{x:\underline{\operatorname{ht}}_{x}I\leq h\}|>\varepsilon_{\beta}n^{2}$, then up to a change of the constant $C$, one has
$1$ $\geq\sum_{I\in\mathbf{I}_{\mathsf{bad}}}\bar{\pi}_{n}^{\mathsf{fl}}(\mathcal{I}_{\mathsf{top}}=\Theta_{H}I)\Big{(}1+e^{-(4\beta+C)h}\Big{)}^{|\{x:\underline{\operatorname{ht}}_{x}I\leq h\}|}\geq\sum_{I\in\mathbf{I}_{\mathsf{bad}}}\bar{\pi}_{n}^{\mathsf{fl}}(\mathcal{I}_{\mathsf{top}}=\Theta_{H}I)\Big{(}1+e^{-(4\beta+C)h}\Big{)}^{\varepsilon_{\beta}n^{2}}\,.$
@@
$1$ $\geq\Big{(}1+e^{-(4\beta+C)h}\Big{)}^{\varepsilon_{\beta}n^{2}}\sum_{I\in\mathbf{I}_{\mathsf{bad}}}e^{-(4\beta+C)Hn}\bar{\pi}_{n}^{\mathsf{fl}}(\mathcal{I}_{\mathsf{top}}=I)$
@@ 2.5. Extending the lower bound to the blue interface
We would now like to extend to $\mathcal{I}_{\mathsf{blue}}$ the lower bound of $\frac{1}{4\beta+C}\log n$ obtained in Proposition 2.20 on the height of most plaquettes in $\mathcal{I}_{\mathsf{top}}$. Note that, to that end, the additive error from Corollary 2.15 (whereby $\mathcal{I}_{\mathsf{blue}}$ is sandwiched between $\mathcal{I}_{\mathsf{bot}},\mathcal{I}_{\mathsf{top}}$) would be far too crude; instead, we will rely on the following result.
@@
with $K>0$ the absolute constant from Proposition 2.3 and $C>0$ another absolute constant. Note that, unlike the more careful analysis of the map $\Phi_{1}$ in Section 2.4.1 (which would correspond to bounding $|\sum_{f\in I}{\mathbf{g}}(f,I)-\sum_{f^{\prime}\in I_{0}}{\mathbf{g}}(f^{\prime},I_{0})|$ via absorbing $K$ for each $f\in I\oplus I_{0}$, and another constant contribution for each such $f$ from summing over the exponential decay of its effect on other plaquettes through the $\mathbf{g}$} term) we resorted to the naive bound $(|I|+|I_{0}|)\|{\mathbf{g}}\|_{\infty}\leq(2+s)Kn^{2}$; this is due to $I_{0}$ being adjacent to the boundary of the box, whereby $r(f,I;f^{\prime},I_{0})$ is $1$ for all $f^{\prime}\in I_{0}$ (demonstrating some of the difficulty in overcoming the potential pinning effect of the hard floor).
@@ 3. Sharp upper bound
    there exists a separating set $\mathcal{S}_{N}$ in $\omega\wedge\omega^{\prime}$ between $\mathcal{L}_{3\ell/4}$ and $\mathcal{L}_{\ell}$, and
@@
On the event $\mathcal{G}$, expose the highest realizing $\mathcal{S}_{N}$ and the lowest realizing $\mathcal{S}_{S}$. Even though there is no order on separating surfaces, the highest (resp., lowest) such is well-defined as the outer boundary of the set generated by exposing the $1$-connected component of $\partial_{N}\Lambda_{n,\ell}^{\prime}$ (resp., $\partial_{S}\Lambda_{n,\ell}^{\prime}$) in the set of plaquettes dual to edges that are closed.
The highest $\mathcal{S}_{N}$ is measurable with respect to the states of edges above it, and the lowest $\mathcal{S}_{S}$ with respect to the states of edges below it. Therefore, by the domain Markov property, we can resample the configurations $(\omega,\sigma)$ and $(\omega^{\prime},\sigma^{\prime})$ in the region sandwiched between these separating surfaces using the identity coupling (the colorings on $\mathcal{S}_{N}$ must be $\mathsf{red}$ in both, and on $\mathcal{S}_{S}$ $\mathsf{blue}$ in both, so indeed the measures induced on the regions sandwiched by the separating surfaces have the same distribution). Since $\mathcal{I}_{\mathsf{full}}$ and therefore $\mathcal{I}_{\mathsf{blue}}$ must be sandwiched between these two separating surfaces when $|\mathsf{h}|\leq\ell/2$, on the event $\mathcal{G}$ we have
@@
The argument for the floor measures $\mu_{\Lambda_{n,\ell}}^{\mathsf{fl}}$ and $\mu_{\Lambda_{n,m}}^{\mathsf{fl}}$ is analogous, only needing the existence of $\mathcal{S}_{N}$, and no need for $\mathcal{S}_{S}$ since the southern boundary is already shared at height zero. The only exception is that we cannot apply Theorem 2.5 directly to bound ${\mathbf{P}}_{\Lambda_{n,\ell}}^{\mathsf{fl}}(\mathcal{I}_{\mathsf{full}}\cap\mathcal{L}_{3\ell/4}\neq\emptyset)$, so we will additionally use a monotonicity argument involving $\mathcal{I}_{\mathsf{blue}}$ (in the same way that we proved an upper bound for $\mathcal{I}_{\mathsf{top}}$ going through $\mathcal{I}_{\mathsf{blue}}$ first in Section 2.2). Compare ${\mathbf{P}}_{\Lambda_{n,\ell}}^{\mathsf{fl}}$ to ${\mathbf{P}}_{\Lambda_{n,\ell}}^{\ell/2}$. The latter is just a shift of the measure ${\mathbf{P}}_{\Lambda^{\prime}_{n,\ell/2}}^{\mathsf{dob}}$ by $\ell/2$, so we have by Theorem 2.5 that
${\mathbf{P}}_{\Lambda_{n,\ell}}^{\ell/2}(\mathcal{I}_{\mathsf{full}}\cap\mathcal{L}_{3\ell/4}\neq\emptyset)={\mathbf{P}}_{\Lambda_{n,\ell/2}^{\prime}}^{\mathsf{dob}}(\mathcal{I}_{\mathsf{full}}\cap\mathcal{L}_{\ell/4}\neq\emptyset)\leq e^{-c\beta n}\,.$
Since $\mathcal{I}_{\mathsf{blue}}\subset\mathcal{I}_{\mathsf{full}}$, the same bound applies for $\mathcal{I}_{\mathsf{blue}}$. Now looking at the Potts marginals, the fuzzy Potts monotonicity and the fact that $\max_{x}\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{blue}})$ is increasing in the set of $\mathsf{blue}$ vertices implies that we also have
$\mu_{\Lambda_{n,\ell}}^{\mathsf{fl}}(\mathcal{I}_{\mathsf{blue}}\cap\mathcal{L}_{3\ell/4}\neq\emptyset)\leq e^{-c\beta n}\,.$
(The same argument is done at the end of the proof of Proposition 2.9, see for details.) Finally, this implies an upper bound of ${\mathbf{P}}_{\Lambda_{n,\ell}}^{\mathsf{fl}}(\mathcal{I}_{\mathsf{full}}\cap\mathcal{L}_{7\ell/8}\neq\emptyset)\leq e^{-c\beta n}$ by Corollary 2.15. ∎